A frog is at the bottom of a 20 foot well. Each day it crawls up 4 feet but each night it slips back 3 feet. How many days are required for the frog to rach the top of the well?
Answer by checkley77(12844)

4-3=1 ft. gain each day.
15 days=15 ft. gained.
The 16th. day 15+4=19-3=16 ft.
The 17th. day 16+4=20 ft. & out for the frog on the 17th. day.